# Do Sell-side Analyst Reports Have Investment Value?

### v1

@@ 1 Introduction
In this paper, I analyze 1.2 million analyst reports from Thomson Reuters Investext between 2000 and 2023 to address two fundamental questions: whether analyst reports contain incremental investment value, and if so, what drives this value? Using this comprehensive dataset, I propose an integrated framework that examines investment value through three lenses: information representation, information relevance, and information interpretation.
@@
I find that the Strategic Outlook discussions, despite comprising only 15% of report content, generate the largest contribution to investment value, accounting for 41% of the portfolio Sharpe ratio. By contrast, Financial Analysis—the largest category by distribution—contributes only about 16% of the Sharpe ratio, suggesting that standard financial data are quickly priced in, leaving less scope for future return prediction. Meanwhile, Company and Industry Overview and Risk and Governance show moderate yet meaningful contributions. Additional Content adds little. Portfolio performance validates the Shapley value attribution analysis. A trading strategy based solely on strategic content generates a monthly return of 1.41% with a Sharpe ratio of 0.93, significantly outperforming other content categories, which yield more modest returns between 0.52% and 0.58% monthly with Sharpe ratios of 0.37 to 0.41.
@@
Strategic Outlook includes both investment thesis and valuation discussions, offering forward-looking elements such as growth potential, valuation, and investment opportunities. For instance, consider this excerpt from a report with a positive future return prediction despite a downgrade: “Although the shares are an attractive long-term investment, we must caution that the severity of the current industry downturn increases the chance of further near-term earnings surprises.” Linking the attribution result with the overreaction story, the investment opportunity appears to arise when analysts identify strong long-term prospects that are temporarily obscured by near-term challenges. Supportively, when dividing the strategic discussion further into three dimensions of timeframe, sentiment, and focus, the Shapley value decomposition emphasizes the contribution of sentences related to positive long-term fundamentals.
@@
A potential concern about return predictability using LLMs is look-ahead bias (Sarkar and Vafa, 2024). Specifically, pre-trained language models could have learned future stock performance information from their training data, potentially causing information leakage through the embedding-generating process. However, I argue the results are robust to this concern for three reasons. First, LLMs are fundamentally designed to optimize language understanding rather than financial prediction capabilities. In other words, the pre-training process focuses on capturing semantic meaning through next-token prediction, not predicting stock returns. Second, if LLMs were indeed learning and exploiting future market information, we would expect to see the strongest predictability with the Company and Industry Overview content, as it contains the most comprehensive stock-specific information. Instead, this category generates insignificant alpha. Third, I directly test for look-ahead bias by examining portfolio performance after each model’s knowledge cutoff date using four different LLMs: BERT, RoBERTa, LLaMA2, and LLaMA3, which were trained on data ending between December 2018 and March 2023. The earliest models (BERT and RoBERTa) provide over five years of true out-of-sample testing. If the predictability stemmed from information leakage, performance should deteriorate in these post-knowledge cutoff periods. Instead, the strategies generate even superior performances, yielding monthly long-short returns between 1.76% and 5.28% with Sharpe ratios ranging from 0.74 to 2.37.
@@
This paper also contributes to the extensive overreaction literature by uncovering a novel pattern in analyst texts: negative sentiment paired with an overall positive long-run outlook.  [^2] For example: De Long et al. (1990); Porta et al. (1997); Hong and Stein (1999); Gennaioli et al. (2016); Weber (2018). A related paper by Bordalo et al. (2019) examines long-term growth expectations (LTG) and finds that analysts overreact to short-term news, with return predictability arising as prices gradually correct these expectations. A major distinction between our findings lies in the asymmetry of overreaction. While the LTG factor is mainly driven by analysts being over-optimistic and overvaluing stocks in high-LTG portfolios, my findings reveal that analysts often exhibit cautionary hedge—through downgrades or negative tones—while simultaneously embedding optimistic forecasts about longer-term fundamentals in their written discussions. The mismatch between near-term pessimism and constructive long-term insights provides a textual signature of market overreaction, where analysts behave cautiously with conservative narratives and forecasts despite being aware of positive fundamentals. By detecting these subtle, dual-layer signals in analyst reports, this work extends prior studies of De Bondt and Thaler (1990) on analyst overreaction and clarifies how overreaction dynamics are expedited with analyst written reports.
Third, this study contributes to the expanding literature on text factors, which has predominantly extracted sentiment or topics from corporate disclosures, earnings-call transcripts, and news media using dictionary-based or topic-modeling approaches (e.g., Tetlock, 2007; Chen et al., 2022; Meursault et al., 2023; Bybee et al., 2023; Hirshleifer et al., 2023). By contrast, I deploy state-of-the-art LLM embeddings to the nuanced, forward-looking narratives in sell-side analyst reports. These embeddings not only capture richer semantic and contextual relations but also have an additive property. The latter allows me to isolate specific content categories—such as financial analysis, risk factors, and strategic outlook—and then quantify each category’s standalone contribution to investment value through Shapley value decomposition. This approach offers a general framework for interpreting granular insights and attributing text predictive power to interested components.
@@ 2.2.3 Shapley Value Decomposition
I implement this framework by mapping tokens to players, topics to teams, and monthly investment performance to individual games. The Shapley value decomposition process aggregates tokens within their respective topics, evaluates each topic’s performance across all possible coalitions in each monthly period, and computes the final contribution as the average marginal contribution of each topic across all periods. Below I present the mathematical formulations.
In implementing the Shapley framework, I first aggregate token embeddings into topic-level representations. LLMs generate token-level embeddings, which I combine into a report-level representation through averaging:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (3)
@@
With outcomes from all possible topic combinations, I can compute the Shapley value using the following formula:
@@ 3.1.1 Return Predictability
To investigate these price dynamics, I use the predicted 12-month ahead stock returns from regression 1 as a measure of analysts’ outlook. I sort reports into decile groups based on their predictions for each trading day. [^4] In the sample, a trading day contains an average of 167 reports. To calculate buy-and-hold abnormal returns, I follow the characteristics-based benchmark approach proposed by Daniel et al. (1997):
@@
where $RET_{it,12m}$ represents the realized 12-month ahead return of stock $i$. $x_{i,t}$ denotes analyst information from each report, consisting of four measures: $\widehat{RET_{12m}}$ is the predicted 12-month ahead return from Ridge regression (1); $REC_{rev}$ represents recommendation revision, calculated as the difference between the current report’s recommendation and the analyst’s last recommendation for the same stock in I/B/E/S; $EF_{rev}$ denotes earnings forecast revision; and $TP_{rev}$ represents target price revision. The model includes firm fixed effects $\alpha_{t}$ to control for systematic market conditions. Standard errors are two-way clustered at both stock and year-quarter levels.
@@ 3.1.2 Short-term Overreaction
Table 4 presents two noteworthy findings. First, the realized returns ($R_{12m}$) exhibit a significant spread of 5%, which closely aligns with the ’H-L’ return spread documented in Figure 1. Second, the contemporaneous market reaction ($CAR_{[0,+1]}$) following report releases shows a negative relationship with predicted returns, with a significant high-minus-low spread of -0.25%. Together with the evidence in Figure 3, these results present an intriguing puzzle: The most optimistic forward-looking reports tend to follow bad news, and vice versa. I propose three potential explanations for this pattern. First, both analysts and the market may overreact to short-term negative news (e.g., a weak quarter or a disappointing press release). Yet the analyst’s deeper analysis (e.g., the model-based 12-month forecast) might conclude that the fundamentals remain solid, eventually driving future price appreciation. Under this scenario, we should observe consistent negative relationships between analyst recommendations, narrative tones, and predicted returns.
@@
Taken together, the data reveal that analysts and the market place disproportionate weight on immediate setbacks, even while the analyst’s expertise anticipates subsequent stock outperformance. In other words, the investment value arises from analysts’ ability to recognize longer-term fundamentals despite their own cautious or negative language in the short run. This systematic mismatch—where near-term negativity coincides with strong future returns—ultimately generates the portfolio alpha.
@@ 3.1.3 Cross-sectional Analyses
I first examine the return predictability of analyst reports across large vs small stocks and mature vs young stocks. Panel (a) shows that reports on large-capitalization stocks generate significantly higher abnormal return spreads compared to small-cap stocks (4.14% vs 1.74%), with the difference particularly pronounced for stocks predicted to have high returns. This pattern suggests that despite the greater analyst coverage and presumed market efficiency in the large-cap segment, skilled analysts can still identify meaningful mispricing opportunities among these stocks. The life cycle results in Panel (b) indicate that mature firms offer greater potential for profitable analyst recommendations compared to young firms. Reports on mature firms predicted to have high returns generate cumulative abnormal returns of approximately 5.15% over the 252-day horizon, while similar reports on young firms yield a substantially lower return spread of 1.51%.
@@
These findings suggest that the investment value of research reports is concentrated among experienced, skilled analysts covering large, mature firms. This pattern is consistent with the idea that superior information production requires both analytical capability and deep institutional knowledge. Moreover, large, mature businesses with more predictable cash flows and stable competitive moats often snap back from negative news—leading to a positive return drift that skilled analysts can exploit.
@@ 3.2 What content matters?
A follow-up question would be, since Shapley value attribution pins down the content contributing the most to investment value, can we separate this most valuable information and form a portfolio free of text noise? In Table 7, I report the value-weighted decile portfolio performance sorted by category-specific content. I focus on four categories, excluding additional content that primarily consists of appendices and latent sentences.
Strategic Outlook content generates the strongest portfolio performance, with the long-short strategy delivering a mean return of 1.41% per month and a Sharpe ratio of 0.93. The outperformance appears to be driven by both the long and short legs, with the highest decile earning 1.87% monthly and showing significant alpha (0.72, t = 2.85). This superior performance aligns with the earlier Shapley value decomposition, which identifies Strategic Outlook as the most valuable content category.
The remaining categories - Company and Industry Overview, Financial Analysis, and Risk and Governance - generate more modest performance metrics, with long-short returns ranging from 0.52% to 0.58% monthly and Sharpe ratios between 0.37 and 0.41. While the highest deciles of these categories show some predictive power, particularly Risk and Governance with an alpha of 0.32 (t = 1.97), their overall performance is substantially weaker than Strategic Outlook. This pattern is consistent with the Shapley value decomposition and suggests that traditional financial analysis, industry insights, and risk assessments are more efficiently priced by the market than forward-looking strategic insights.
Therefore, isolating Strategic Outlook content may indeed help form more efficient portfolios by focusing on the most valuable component of analyst research. Confirming the effectiveness of the attribution analysis with Shapley values, the results again highlight the long-term investment opportunities embedded in analysts’ forward-looking strategic insights.
@@ 3.3 Incremental Investment Value
Given the ”zoo of factors” documented to have investment value, a key question emerges: is the investment value of analyst reports incremental to existing predictive signals from analyst forecasts and fundamental factors? I therefore extend my analysis to examine whether analyst reports contain unique information beyond two comprehensive sets of established predictors: 18 analyst-based factors including EPS forecast revisions and recommendation changes from the Chen and Zimmermann (2022) database, and 94 market and fundamental-based factors documented to have cross-sectional predictability in Gu et al. (2020). [^5] The open-source code is publicly available on Dacheng Xiu’s website. Detailed descriptions of these analyst-based and fundamental-based factors can be found in Tables A2 and A1, respectively.
@@
Panel A first establishes the standalone performance of each information source. Both full report content (RP) and strategic outlook sections (SO) generate significant risk-adjusted returns across various factor models, with alphas ranging from 0.73 to 1.21 for RP and 0.95 to 1.46 for SO. The traditional analyst-based factors (ANA) show modest but significant predictive power (alphas between 0.16 and 0.28). The machine learning-based anomaly portfolio (ANOM) exhibits strong performance with alphas between 1.15 and 1.40, suggesting effective capture of fundamental signals.
Panels B and C address the central question of incremental value. When combining report predictions with analyst-based factors (RP + ANA), the portfolio maintains significant alphas (0.50-0.68) and shows meaningful information gain with an IR of 0.73. The combination with fundamental factors (RP + ANOM) yields even stronger results, with alphas between 1.03 and 1.18 and an IR of 1.16. Most notably, incorporating both analyst and fundamental factors (RP + ANA + ANOM) produces the highest Sharpe ratio (1.60) and consistent alphas across factor models (0.77-0.92), with an IR of 1.23 (t=4.05). The patterns are similar but even stronger when using strategic outlook content instead of full reports (IR=1.41, t=4.35), consistent with my earlier findings about the particular value of strategic discussions.
@@ 3.4 Look ahead Bias
A key concern in demonstrating out-of-sample profitability from analyst reports is look-ahead bias. Because LLMs are pre-trained on massive text corpora, there is a risk that they could have indirectly learned from future stock performance data. While Table 7 shows that providing corporate and industry information to LLMs does not generate significant profits—mitigating concerns about entity- or industry-level information leakage—one might argue that strategic outlooks contain forward-looking information that could align with the LLMs’ pre-training corpus. The fact that these outlooks can later be confirmed by actual outcomes may introduce a potential source of information leakage.
To address this concern, I conduct a stringent analysis restricted to post-knowledge cutoff samples. This analysis encompasses a broad range of language models with diverse architectures and scales. BERT and RoBERTa are relatively smaller, pioneering transformer-based NLP models (with 110 million to 355 million parameters). In contrast, LLaMA2-13B is a significantly larger model at 13 billion parameters, while LLaMA3-8B represents a more recent architecture with 8 billion parameters. Specifically, I evaluate portfolio performance only after each model’s training data cutoff date: BERT (post-June 2024), RoBERTa (post-June 2024), LLaMA2 (post-June 2024), and LLaMA3 (post-April 2024). Table 9 presents these results across two dimensions: analysis of full reports (Panel A) and strategic outlooks (Panel B).
Strikingly, the performance in this true out-of-sample period is even stronger than in the full sample. Examining the high-minus-low (H-L) portfolio returns in Panel A, all models generate substantial positive returns ranging from 1.59% (LLaMA2) to 2.70% (LLaMA3) with corresponding Sharpe ratios between 0.74 and 1.28. The performance amplifies further when focusing solely on strategic outlooks (Panel B), with H-L returns more than doubling for some models, ranging from 1.81% (RoBERTa) to an impressive 5.28% (LLaMA3). The enhanced performance in the post-cutoff period provides compelling evidence that the LLMs’ ability to extract profitable information from analyst reports is not driven by potential information leakage from their training data.

### v2

@@ 1 Introduction
In this paper, I analyze 1.2 million analyst reports from Thomson Reuters Investext between 2000 and 2023 to address two fundamental questions: whether analyst reports contain incremental investment value and, if so, what drives this value. Using this comprehensive dataset, I propose an integrated framework that examines investment value through three lenses: information representation, information relevance, and information interpretation.
@@
I find that the Strategic Outlook discussions, despite comprising only 15% of report content, generate the largest contribution to investment value, accounting for 41% of the portfolio Sharpe ratio. By contrast, Financial Analysis—the largest category by distribution—contributes only about 16% of the Sharpe ratio, suggesting that standard financial data are quickly priced in, leaving less scope for future return prediction. Meanwhile, Company and Industry Overview and Risk and Governance show moderate contributions. Additional Content adds little. Univariate portfolio performance validates the Shapley value attribution analysis. A trading strategy based solely on strategic content generates a monthly long-short return of 1.41% with a Sharpe ratio of 0.93, significantly outperforming other content categories, which yield more modest returns between 0.52% and 0.58% monthly with Sharpe ratios of 0.37 to 0.41.
@@
Strategic Outlook includes forward-looking elements such as growth potential, valuation, and investment opportunities. Consider this excerpt from a report with a positive future return prediction despite a downgrade: “Although the shares are an attractive long-term investment, we must caution that the severity of the current industry downturn increases the chance of further near-term earnings surprises.” Linking the attribution result with the overreaction story, the investment opportunity appears to arise when analysts identify strong long-term prospects that are temporarily obscured by near-term challenges. Supportively, when dividing the strategic discussion further into three dimensions of timeframe, sentiment, and focus, the Shapley value decomposition emphasizes the contribution of sentences related to positive long-term fundamentals.
@@
A potential concern about return predictability is look-ahead bias with LLMs (Sarkar and Vafa, 2024). Since pre-trained language models could have learned future stock performance information from their training data, the embedding-generating process could introduce information leakage. However, I argue that the results are robust to this concern for three reasons. First, LLMs are fundamentally designed to optimize language understanding rather than financial prediction capabilities. In other words, the pre-training process focuses on capturing semantic meaning through next-token prediction, not predicting stock returns. Second, if LLMs were indeed learning and exploiting future market information, we would expect to see the strongest predictability with the Company and Industry Overview content, as it contains the most comprehensive stock-specific information. Instead, this category generates insignificant alpha. Third, I directly test for look-ahead bias by examining portfolio performance after each model’s knowledge cutoff date using four different LLMs: BERT, RoBERTa, LLaMA2, and LLaMA3, which were trained on data ending between December 2018 and March 2023. The earliest models (BERT and RoBERTa) provide over five years of true out-of-sample testing. If the predictability stemmed from information leakage, performance should deteriorate in these post-knowledge cutoff periods. Instead, the strategies generate even superior performances, yielding monthly long-short returns between 1.76% and 5.28% with Sharpe ratios ranging from 0.74 to 2.37.
@@
This paper also contributes to the extensive overreaction literature by uncovering a novel pattern in analyst texts: negative sentiment paired with an overall positive long-run outlook. [^2] For example: De Long et al. (1990); Porta et al. (1997); Hong and Stein (1999); Gennaioli et al. (2016); Weber (2018). A related paper by Bordalo et al. (2019) examines long-term growth expectations (LTG) and finds that analysts overreact to short-term news, with return predictability arising as prices gradually correct these expectations. A major distinction between our findings lies in the asymmetry of overreaction. While the LTG factor is mainly driven by analysts being over-optimistic and overvaluing stocks in high-LTG portfolios, my findings reveal that analysts often exhibit cautionary hedge through downgrades or negative tones, while simultaneously embedding optimistic forecasts about longer-term fundamentals in their written discussions. The mismatch between near-term pessimism and constructive long-term insights provides a textual signature of market overreaction, where analysts behave cautiously with conservative narratives and forecasts despite being aware of positive fundamentals. By detecting these subtle, dual-layer signals in analyst reports, this work extends prior studies of De Bondt and Thaler (1990) on analyst overreaction and clarifies how overreaction dynamics are expedited with analyst written reports.
Third, this study contributes to the expanding literature on text factors, which has predominantly extracted sentiment or topics from corporate disclosures, earnings-call transcripts, and news media using dictionary-based or topic-modeling approaches (e.g., Tetlock, 2007; Chen et al., 2022; Meursault et al., 2023; Bybee et al., 2023; Hirshleifer et al., 2023). By contrast, I deploy state-of-the-art LLM embeddings to the nuanced, forward-looking narratives in sell-side analyst reports. These embeddings not only capture richer semantic and contextual relations but also have an additive property. The latter allows me to isolate specific content categories, such as financial analysis, risk factors, and strategic outlook, and then quantify each category’s standalone contribution to investment value through Shapley value decomposition. This approach offers a general framework for interpreting granular insights and attributing text predictive power to interested components.
@@ 2.2.3 Shapley Value Decomposition
I implement this framework by mapping tokens to players, topics to teams, and monthly investment performance to individual games. The Shapley value decomposition process aggregates tokens within their respective topics, evaluates each topic’s performance across all possible coalitions in each monthly period, and computes the final contribution as the average marginal contribution of each topic across all periods. Below, I present the mathematical formulations.
I first aggregate token embeddings into topic-level representations. LLMs generate token-level embeddings, which I combine into a report-level representation through averaging:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}{e}_{i},$ (3)
@@
With outcomes from all possible topic combinations, I then compute the Shapley value using the following formula:
@@ 3.1.1 Return Predictability
To investigate these price dynamics, I use the predicted 12-month ahead stock returns from regression (1) as a measure of analysts’ outlook. I sort reports into decile groups based on their predictions for each trading day. [^4] In the sample, a trading day contains an average of 167 reports. To calculate buy-and-hold abnormal returns, I follow the characteristics-based benchmark approach proposed by Daniel et al. (1997):
@@
where $RET_{it,12m}$ represents the realized 12-month ahead return of stock $i$. $x_{i,t}$ denotes analyst information from each report, consisting of four measures: $\widehat{RET_{12m}}$ is the predicted 12-month ahead return from Ridge regression (1); $REC_{rev}$ represents recommendation revision, calculated as the difference between the current report’s recommendation and the analyst’s last recommendation for the same stock in I/B/E/S; $EF_{rev}$ denotes earnings forecast revision; and $TP_{rev}$ represents target price revision. The model includes year-month fixed effects $\alpha_{t}$ to control for systematic market conditions. Standard errors are clustered two ways at both stock and year-month levels.
@@ 3.1.2 Short-term Overreaction
Table 4 presents two noteworthy findings. First, the realized returns ($R_{12m}$) exhibit a significant spread of 5%, which closely aligns with the ’H-L’ return spread documented in Figure 1. Second, the contemporaneous market reaction ($CAR_{[0,+1]}$) following report releases shows a negative relationship with predicted returns, with a significant high-minus-low spread of -0.25%. Together with the evidence in Figure 3, these results present an intriguing puzzle: The most optimistic forward-looking reports tend to follow bad news, and vice versa.
I propose three potential explanations for this pattern. First, both analysts and the market may overreact to short-term negative news (e.g., a weak quarter or a disappointing press release). Yet the analyst’s deeper analysis (e.g., the model-based 12-month forecast) might conclude that the fundamentals remain solid, eventually driving future price appreciation. Under this scenario, we should observe consistent negative relationships between analyst recommendations, narrative tones, and predicted returns.
@@
Taken together, the analyses reveal that analysts and the market place disproportionate weight on immediate setbacks, even while the analyst’s expertise anticipates subsequent stock outperformance. The investment value arises from analysts’ ability to recognize longer-term fundamentals despite their own cautious or negative language in the short run. This systematic mismatch, where near-term negativity coincides with strong future returns, ultimately generates the portfolio alpha.
@@ 3.1.3 Cross-sectional Analyses
I first examine the return predictability of analyst reports across large vs small stocks and mature vs young stocks. Panel (a) shows that reports on large-capitalization stocks generate significantly higher abnormal return spreads compared to small-cap stocks (4.14% vs 1.74%), with the difference particularly pronounced for stocks predicted to have high returns. This pattern suggests that despite the greater analyst coverage and presumed market efficiency in the large-cap segment, reports from skilled analysts still contain meaningful investment opportunities among these stocks. The life cycle results in Panel (b) indicate that mature firms offer greater potential for profitable analyst recommendations compared to young firms. Reports on mature firms predicted to have high returns generate cumulative abnormal returns of approximately 5.15% over the 252-day horizon, while similar reports on young firms yield a substantially lower return spread of 1.51%.
@@
These findings suggest that the investment value of research reports is concentrated among experienced, skilled analysts covering large, mature firms. This pattern is consistent with the idea that superior information production requires both analytical capability and deep institutional knowledge. Moreover, large, mature businesses with more predictable cash flows and stable competitive moats often snap back from negative news—leading to a positive return drift that skilled analysts may hint.
@@ 3.2 What content matters?
A follow-up question would be, since Shapley value attribution pins down the content contributing the most to investment value, can we separate this most valuable information and form a portfolio free of text noise? In Table 7, I report the value-weighted decile portfolio performance sorted by category-specific content. I focus on four categories, excluding additional content that primarily consists of latent sentences.
Strategic Outlook content generates the strongest portfolio performance, with the long-short strategy delivering a mean return of 1.41% per month and a Sharpe ratio of 0.93. The outperformance appears to be driven by both the long and short legs, with the highest decile earning 1.87% monthly and showing significant alpha (0.72, t = 2.85). This superior performance aligns with the earlier Shapley value decomposition, which identifies Strategic Outlook as the most valuable category.
The remaining categories, Company and Industry Overview, Financial Analysis, and Risk and Governance, generate more modest performance metrics. Long-short returns range from 0.52% to 0.58% monthly and Sharpe ratios are between 0.37 and 0.41. While the highest deciles of these categories show some predictive power, particularly Risk and Governance with an alpha of 0.32 (t = 1.97), their overall performance is substantially weaker than Strategic Outlook. This pattern is again consistent with the Shapley value decomposition. It suggests that traditional financial analysis, industry insights, and risk assessments are less informative about future returns than forward-looking strategic insights.
In general, the results again highlight the long-term investment opportunities embedded in analysts’ forward-looking strategic insights. Additionally, the univariate analysis confirms the effectiveness of the attribution analysis with Shapley values, showing that focusing on the most valuable component of Strategic Outlook content indeed helps form more efficient portfolios.
@@ 3.3 Incremental Investment Value
Literature has documented “zoo of factors” with investment value. A key question is whether the investment value of analyst reports is incremental to existing predictive signals. I therefore extend my analysis to examine whether analyst reports contain unique information beyond two comprehensive sets of established predictors: 18 analyst-based factors including EPS forecast revisions and recommendation changes from the Chen and Zimmermann (2022) database, and 94 market and fundamental-based factors documented to have cross-sectional predictability in Gu et al. (2020). [^5] The replication code is publicly available on Dacheng Xiu’s website. Detailed descriptions of these analyst-based and fundamental-based factors can be found in Tables A2 and A1, respectively.
@@
Panel A first establishes the standalone performance of each information source. Both full report content (RP) and strategic outlook sections (SO) generate significant risk-adjusted returns across various factor models, with alphas ranging from 0.73 to 1.21 for RP and 0.95 to 1.46 for SO. The numerical analyst-based factors (ANA) show modest but significant predictive power (alphas between 0.16 and 0.28). The machine learning-based anomaly portfolio (ANOM) exhibits strong performance with alphas between 1.15 and 1.40, suggesting effective capture of fundamental signals.
Panels B and C address the central question of incremental value. When combining report predictions with analyst-based factors (RP + ANA), the portfolio maintains significant alphas (0.50-0.68) and shows meaningful incremental gain with an IR of 0.73. The combination with fundamental factors (RP + ANOM) yields even stronger results, with alphas between 1.03 and 1.18 and an IR of 1.16. Most notably, incorporating both analyst and fundamental factors (RP + ANA + ANOM) produces the highest Sharpe ratio (1.60) and consistent alphas across factor models (0.77-0.92), with an IR of 1.23 (t=4.05). The patterns are similar but even stronger when using strategic outlook content instead of full reports (IR=1.41, t=4.35), consistent with my earlier findings about the particular value of analyst strategic outlook discussions.
@@ 3.4 Look ahead Bias
A key concern in demonstrating out-of-sample profitability from analyst reports is look-ahead bias. Because LLMs are pre-trained on massive text corpora, there is a risk that they could have indirectly learned from future stock performance data. While Table 7 shows that providing corporate and industry information to LLMs does not generate significant profits—mitigating concerns about entity- or industry-level information leakage—one might argue that strategic outlooks contain forward-looking information that could align with the LLMs’ pre-training corpus. The fact that these outlooks can later be confirmed by actual outcomes may introduce another source of information leakage.
To further address this concern, I conduct a stringent analysis restricted to post-knowledge cutoff samples. This analysis encompasses a broad range of language models with diverse architectures and scales. BERT and RoBERTa are relatively smaller, pioneering transformer-based NLP models (with 110 million to 355 million parameters). In contrast, LLaMA2-13B is a significantly larger model at 13 billion parameters, while LLaMA3-8B represents a more recent architecture with 8 billion parameters. Specifically, I evaluate portfolio performance only after each model’s training data cutoff date: BERT (post-June 2024), RoBERTa (post-June 2024), LLaMA2 (post-June 2024), and LLaMA3 (post-April 2024). Table 9 presents these results across two dimensions: analysis of full reports (Panel A) and strategic outlooks (Panel B).
Strikingly, the performance in this true out-of-sample period is even stronger than in the full sample. Examining the high-minus-low (H-L) portfolio returns in Panel A, all models generate substantial positive returns ranging from 1.59% (LLaMA2) to 2.70% (LLaMA3) with corresponding Sharpe ratios between 0.74 and 1.28. The performance amplifies further when focusing solely on strategic outlooks (Panel B), with H-L returns more than doubling for some models, ranging from 1.81% (RoBERTa) to an impressive 5.28% (LLaMA3). The enhanced performance in the post-cutoff period provides compelling evidence that the LLMs’ ability to extract profitable information from analyst reports is not driven by potential information leakage from their training text.